OpenStudy (anonymous):

Solve the triangle. A = 52°, b = 14, c = 6

OpenStudy (ajprincess):

R u familiar with cos rule?

OpenStudy (anonymous):

Yes, I already have side a as 11.3 but I need help finding the other two angles B and C!

OpenStudy (ajprincess):

u can use sine rule

OpenStudy (anonymous):

I did and I got 77.5 degrees for B and that's not in my answer options :(

OpenStudy (ajprincess):

what r ur answr options?

OpenStudy (anonymous):

a ≈ 14.9, C ≈ 24.2, B ≈ 103.8 a ≈ 11.3, C ≈ 24.2, B ≈ 103.8 a ≈ 14.9, C ≈ 28.2, B ≈ 99.8 No triangles possible so far I have the 3rd one as my answer but i still want to make sure C=24.2 and B=103.8 if that's even the right answer,

OpenStudy (anonymous):

heellp meeee!

OpenStudy (raden):

i will go the 2nd one

OpenStudy (raden):

its depend on digits decimal, i got B = 76.704 use the identity : sinx = sin ( 180 - x) so, sin76.704 = sin(180 - 76.704) = sin103.296 it nearest B = 103.8

OpenStudy (raden):

A+B+C = 180 C = 180 - (A+B) C = 180 - (52 + 103.8) C = 180 - 155.8 C = 24.2

OpenStudy (anonymous):

Thank you !